# AdvLoRA: Adversarial Low-Rank Adaptation of Vision-Language Models

### v1

@@ 3.3. Adaptation Module
Although adversarial training can effectively enhance VLMs’ adversarial robustness, it requires updating all parameters based on gradient information, leading to a significant cost overhead. To alleviate this issue, in this subsection, we propose an adaptation module that performs adversarial training on LoRA to reduce the number of learnable parameters, achieving parameter-efficient adversarial adaptation. We first provide a brief introduction to LoRA, followed by the introduction of clustering reparameterization and parameter alignment methods, as well as an adaptive parameter update strategy, to facilitate adversarial adaptation.
3.3.1. LoRA
LoRA achieves parameter-efficient adaptation by updating two low-rank matrices attached to the frozen pre-trained weights. Specifically, given the pre-trained weights $\mathbf{W_{0}}\in\mathbb{R}^{m\times n}$, and the LoRA matrices $\mathbf{A}\in\mathbb{R}^{m\times k}$, $\mathbf{B}\in\mathbb{R}^{k\times n}$, the input $\mathbf{X}^{(l-1)}\in\mathbb{R}^{b\times m}$ is processed through the following computation to obtain the output $\mathbf{X}^{(l)}\in\mathbb{R}^{b\times n}$ as follows.
(5) $\mathbf{X}^{(l)}=\mathbf{X}^{(l-1)}\mathbf{W_{0}}+\mathbf{X}^{(l-1)}\mathbf{AB},$
where $k\ll\min(m,n)$. And $\mathbf{A}$ and $\mathbf{B}$ are initialized as follows.
(6) $\mathbf{A}\sim\mathcal{N}(0,\sigma^{2}),\quad\mathbf{B}=\textbf{0},$
@@ 3.3.1. LoRA
During the adaptation process, $\mathbf{W_{0}}$ is fixed, while $\mathbf{A}$ and $\mathbf{B}$ are updated via the gradient descent methods. In our proposed model, AdvLoRA, we freeze $\mathbf{W_{0}}$ and solely update $\mathbf{A},\mathbf{B}$ through adversarial adaptation to achieve adversarial robustness in the model as follows.
(7) $\theta_{\mathbf{A},\mathbf{B}}=\mathop{\arg\min}_{\theta_{\mathbf{A},\mathbf{B}}}\mathcal{L}(v_{a},w).$
@@ 3.3.2. Reparameterization and Adaptive Parameter Update
The primary distinction between AdvLoRA and other LoRA-like methods lies in the parameterization process of the matrices $\mathbf{A},\mathbf{B}$. In the original LoRA, a random Gaussian initialization for $\mathbf{A}$ and zero for $\mathbf{B}$, so $\mathbf{AB}$ is zero at the beginning of adaptation. In contrast, our model, AdvLoRA, initially performs clustering on the weight matrix $\mathbf{W_{0}}$ of the pre-trained model, treating the rank $k$ of LoRA as the number of cluster centers. Specifically, given an weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ and the rank $k$, we first randomly initialize $k$ cluster center $\bm{C}=\{\mathbf{c_{1}},\mathbf{c_{2}},\ldots,\mathbf{c_{k}}\}$. Then, for each column $\mathbf{w_{i}}$ of $\mathbf{W}$, compute the distances to each cluster center $\mathbf{c_{j}}$ and assign $\mathbf{w_{i}}$ to the closest cluster as follows.
(8) $\textrm{cluster}_{i}=\mathop{\arg\min}_{j}\quad\|{\mathbf{w_{i}}-\mathbf{c_{j}}}\|_{2}.$
@@
Then update the cluster centers by computing the mean of all data points assigned to each cluster as follows.
(9) $\mathbf{c_{j}}=\frac{1}{|\mathbf{S}_{j}|}\sum_{\mathbf{w_{i}}\in\mathbf{S}_{j}}\mathbf{w_{i}},$
@@
where $\mathbf{S}_{j}$ is the set of columns of $\mathbf{W}$ assigned to cluster $j$. Repeat the above steps until the cluster centers no longer change significantly or a maximum number of iterations is reached. In this manner, we obtain the cluster center embeddings $\mathbf{C}\in\mathbb{R}^{k\times n}$ and the distance assignment matrix $\mathbf{D}\in\mathbb{R}^{m\times k}$, where each element $d_{ij}$ represents the distance between the $\mathbf{w}_{i}$ and cluster center $\mathbf{c}_{j}$. The distance assignment matrix $\mathbf{D}$ can be computed using the following formula.
(10) $\mathbf{d}_{ij}=\|\mathbf{w_{i}}-\mathbf{c_{j}}\|_{2}.$
@@
And the cluster center representation matrix $\mathbf{C}$ is simply the matrix of cluster centers as follows.
(11) $\mathbf{C}=\begin{bmatrix}\mathbf{c_{1}},\mathbf{c_{2}},\ldots,\mathbf{c_{k}}\end{bmatrix}.$
@@
After the parameter clustering, the clustering assignment matrix $\mathbf{D}\in\mathbb{R}^{m\times k}$ and the parameter center $\mathbf{C}\in\mathbb{R}^{k\times n}$ can be represented the $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{k\times n}$ in the original LoRA method. By these settings, we provide a better reparameterization of the tunable parameters in LoRA. It separates the parameters into different clusters, which have different functions in the whole network.
@@
After obtaining the matrices $\mathbf{A}$ and $\mathbf{B}$, we further impose constraints on their product $\mathbf{AB}$ to align with the parameter distribution of the original weight matrix $\mathbf{W}_{0}$ as follows.
(12) $\min{\quad\|{\mathbf{W_{0}}-\mathbf{AB}}\|_{2}}.$
@@
In this manner, we can guarantee the zero initialization of $\mathbf{AB}$ at the beginning of the training.
@@
During the process of model adversarial adaptation, we design an adaptive update parameter, $\alpha$, to facilitate the model’s adaptive learning of robust semantic representations as follows.
(13) $\mathbf{Y}=\mathbf{XW_{0}}+\alpha\cdot\mathbf{XAB}.$
@@
$\alpha$ is a learnable neural network parameter, which can control the adaptation rate during the adversarial adaptation. In summary, we delineate the entire workflow of AdvLoRA in Algorithm 1.
@@
0:
Table 1. Vulnerability experiment on MSCOCO. “FFT” and “LP” denoting full fine-tuning and linear probe. “Nat” and “Att” donate natural images and adversarially attacked images. “TR” and “IR” donate text-to-image retrieval and image-to-text retrieval.
Method Tunable Para. MSCOCO TR MSCOCO IR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT+Nat 223M 80.46 95.40 97.64 91.17 63.25 85.54 91.49 80.09 85.63
BLIP+FFT+Att 53.38 75.12 82.62 70.37 42.25 67.03 76.47 61.92 66.15(-19.48%)
BLIP+LP+Nat 0.5M 72.30 91.10 95.22 86.21 56.96 80.75 87.85 75.19 80.70
BLIP+LP+Att 43.22 65.82 74.46 61.17 34.60 58.59 68.86 54.12 57.65(-23.05%)
BLIP+LoRA+Nat 2.8M 70.50 90.28 94.58 85.12 56.39 80.36 87.45 74.73 79.93
BLIP+LoRA+Att 43.20 66.20 74.80 61.40 35.85 60.40 70.16 55.47 58.44(-21.49%)
UniAdapter+Nat 19.5M 79.60 94.50 97.26 90.45 62.53 84.95 90.97 79.49 84.97
UniAdapter+Att 53.98 75.66 82.74 70.79 42.02 66.80 76.39 61.74 66.27(-18.70%)
Aurora+Nat 0.3M 78.00 93.40 96.66 89.35 61.45 83.95 90.39 78.60 83.98
Aurora+Att 44.56 67.04 75.00 62.20 34.98 59.34 68.75 54.36 58.28(-25.69%)
Table 2. Vulnerability experiment on MSR-VTT. “TR” and “VR” donate text-to-video retrieval and video-to-text retrieval, respectively.
Method Tunable Para. MSR-VTT TR MSR-VTT VR
R@1 R@5 R@10 R@Mean R@1 R@5 R@10 R@Mean Mean
BLIP+FFT+Nat 223M 20.3 41.3 53.8 38.47 23.4 48.4 60.8 44.2 41.33
BLIP+FFT+Att 1.2 5 7.6 4.6 2.7 8.1 12.5 7.77 6.18(-35.15%)
BLIP+LP+Nat 0.5M 40.3 63.2 72.0 58.5 41.8 63.7 71.6 59.03 58.77
BLIP+LP+Att 7.7 16.1 20.1 14.63 14.4 26.4 32.8 24.53 19.58(-39.19%)
BLIP+LoRA+Nat 2.8M 47.2 71.4 80.5 66.36 45.8 70.7 80.3 65.6 65.98
BLIP+LoRA+Att 12.8 23.4 28.1 21.43 18.9 3.8 37.8 29.16 25.30(-40.68%)
UniAdapter+Nat 19.5M 42.4 68.4 77.4 62.73 42.9 68.4 78.3 63.2 62.97
UniAdapter+Att 8.3 15.4 18.9 14.2 11.6 22.6 27.2 20.47 17.33(-45.64%)
Aurora+Nat 0.3M 45.1 69.7 79.4 64.73 44.2 68.5 77.8 63.5 64.12
Aurora+Att 11.6 20.3 24.6 18.83 16.9 30.1 36.7 27.9 23.37(-40.75%)
@@ 4.1. Experimental Setup
4.1.1. Datasets
We comprehensively evaluated our proposed model, AdvLoRA, on two types of retrieval tasks and four commonly used datasets, to demonstrate the superior performance of AdvLoRA on cross-modal understanding tasks, including image-text retrieval: Flickr30K (Plummer et al., 2015) and MSCOCO (Lin et al., 2014); as well as video-text retrieval: DiDeMo (Anne Hendricks et al., 2017) and MSRVTT (Xu et al., 2016), More details can be seen in Appendix.
4.1.2. Baselines
We compare AdvLoRA with conventional adaptation methods, which are implemented by BLIP: full fine-tuning (BLIP-FFT), linear probe (BLIP-LP); as well as the PEFT method on BLIP: LoRA(BLIP-LoRA), Aurora, and Uniadapter. See more details in the Appendix.
4.1.3. Metrics
We employ $Recall@k$ as our evaluation metric, where $k$ denotes the number of entries considered within the top $k$ retrieval results. This metric is expressed as a percentage.
4.1.4. Implementations
Our implementation is based on Salesforce’s open-source codebase (Li et al., 2022). Following (Lu et al., 2023; Wang et al., 2023), we also apply BLIP (Li et al., 2022) as our vision-language backbone for all tasks. We use PyTorch to implement all experiments on the NVIDIA V100 GPU (32G). For the video-text retrieval task, we follow the work of Wei et al. (Wei et al., 2019) by adopting an attack strategy that sparsely samples 20% of the video frames. Furthermore, we adopt the setup of BLIP, utilizing a momentum encoder to enhance the retrieval performance of our model. To ensure a fair comparison, the momentum encoder is also applied to the other baseline methods. We use AdamW (Loshchilov and Hutter, 2018) optimizer with weight decay. The rank of our proposed AdvLoRA is 10. Note that during the fine-tuning process, the parameters of the backbone model are kept frozen. More training details can be seen in the Appendix.
4.2. Vulnerability to Adversarial Attacks
In this section, we conduct adversarial attacks on BLIP and their variants adapted using PEFT methods to investigate their vulnerability to such attacks. Specifically, we perform PGD-3 attacks on the baseline model for two tasks across four datasets and then evaluate their performance under adversarial attacks. Figure 1 provides a simple illustration of the models’ vulnerability to adversarial attacks, while Table 1 and Table 2 present detailed data. The complete results on other datasets are provided in the Appendix. Through extensive experimentation, we draw a key conclusion as follows.
BLIP adapted by different methods are highly susceptible to adversarial perturbations. As Table 1 and Table 2 indicate, regardless of whether the method used is full fine-tuning or PEFT, performance degradation of 30.98% is observed. This phenomenon can be attributed to the inability of conventional VLMs and adaptation techniques to effectively learn semantically invariant features from the data.
@@ 4.2. Vulnerability to Adversarial Attacks
Table 4. Adversarial experiment on Flikcr30K. An asterisk (*) indicates that adversarial adaptation has been performed. The best results are displayed in bold, while the second-best results are underlined.

### v3

@@ 3.3. Adaptation Module
Although adversarial training can effectively enhance VLMs’ adversarial robustness, it requires updating all parameters based on gradient information, leading to a significant cost overhead. To alleviate this issue, in this subsection, we propose an adaptation module that performs adversarial training on LoRA to reduce the number of tunable parameters, achieving parameter-efficient adversarial adaptation. We first provide a brief introduction to LoRA, followed by the introduction of clustering reparameterization and parameter alignment methods, as well as an adaptive parameter update strategy, to facilitate adversarial adaptation.
LoRA. LoRA achieves parameter-efficient adaptation by updating two low-rank matrices attached to the frozen pre-trained weights. Specifically, given the pre-trained weights $\boldsymbol{W_{0}}\in\mathbb{R}^{m\times n}$, and the LoRA matrices $\boldsymbol{A}\in\mathbb{R}^{m\times k}$, $\boldsymbol{B}\in\mathbb{R}^{k\times n}$, the input $\boldsymbol{X}^{(l-1)}\in\mathbb{R}^{b\times m}$ is processed through the following computation to obtain the output $\boldsymbol{X}^{(l)}\in\mathbb{R}^{b\times n}$ as follows:
(5) $\boldsymbol{X}^{(l)}=\boldsymbol{X}^{(l-1)}\boldsymbol{W_{0}}+\boldsymbol{X}^{(l-1)}\boldsymbol{AB},$
where $k\ll\min(m,n)$. And $\boldsymbol{A}$ and $\boldsymbol{B}$ are initialized as follows:
(6) $\boldsymbol{A}\sim\mathcal{N}(0,\sigma^{2}),\quad\boldsymbol{B}=\textbf{0},$
@@
During the adaptation process, $\boldsymbol{W_{0}}$ is fixed, while $\boldsymbol{A}$ and $\boldsymbol{B}$ are updated via the gradient descent methods. In our proposed model, AdvLoRA, we freeze $\boldsymbol{W_{0}}$ and solely update $\boldsymbol{A},\boldsymbol{B}$ through adversarial adaptation to achieve adversarial robustness in the model as follows:
(7) $\theta_{\boldsymbol{A},\boldsymbol{B}}=\mathop{\arg\min}_{\theta_{\boldsymbol{A},\boldsymbol{B}}}\mathcal{L}(v_{a},w).$
@@
Reparameterization and Adaptive Parameter Update. The primary distinction between AdvLoRA and other LoRA-like methods lies in the parameterization process of the matrices $\boldsymbol{A},\boldsymbol{B}$. In the original LoRA, a random Gaussian initialization for $\boldsymbol{A}$ and zero for $\boldsymbol{B}$, so $\boldsymbol{AB}$ is zero at the beginning of adaptation. In contrast, our model, AdvLoRA, initially performs clustering on the weight matrix $\boldsymbol{W_{0}}$ of the pre-trained model, treating the rank $k$ of LoRA as the number of cluster centers. Specifically, given an weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$ and the rank $k$, we first randomly initialize $k$ cluster center $\boldsymbol{C}=\{\boldsymbol{c_{1}},\boldsymbol{c_{2}},\ldots,\boldsymbol{c_{k}}\}$. Then, for each column $\boldsymbol{w_{i}}$ of $\boldsymbol{W}$, compute the distances to each cluster center $\boldsymbol{c_{j}}$ and assign $\boldsymbol{w_{i}}$ to the closest cluster as follows:
(8) $\textrm{cluster}_{i}=\mathop{\arg\min}_{j}\|{\boldsymbol{w_{i}}-\boldsymbol{c_{j}}}\|_{2}.$
@@
Then update the cluster centers by computing the mean of all data points assigned to each cluster as follows:
(9) $\boldsymbol{c_{j}}=\frac{1}{|\boldsymbol{S}_{j}|}\sum_{\boldsymbol{w_{i}}\in\boldsymbol{S}_{j}}\boldsymbol{w_{i}},$
@@
where $\boldsymbol{S}_{j}$ is the set of columns of $\boldsymbol{W}$ assigned to cluster $j$. Repeat the above steps until the cluster centers no longer change significantly or a maximum number of iterations is reached. In this manner, we obtain the cluster center embeddings $\boldsymbol{C}\in\mathbb{R}^{k\times n}$ and the distance assignment matrix $\boldsymbol{D}\in\mathbb{R}^{m\times k}$, where each element $d_{ij}$ represents the distance between the $\boldsymbol{w}_{i}$ and cluster center $\boldsymbol{c}_{j}$. The distance assignment matrix $\boldsymbol{D}$ can be computed using the following formula:
(10) $\boldsymbol{d}_{ij}=\|\boldsymbol{w_{i}}-\boldsymbol{c_{j}}\|_{2}.$
@@
And the cluster center representation matrix $\boldsymbol{C}$ is simply the matrix of cluster centers as follows:
(11) $\boldsymbol{C}=\begin{bmatrix}\boldsymbol{c_{1}},\boldsymbol{c_{2}},\ldots,\boldsymbol{c_{k}}\end{bmatrix}.$
@@
After the parameter clustering, the clustering assignment matrix $\boldsymbol{D}\in\mathbb{R}^{m\times k}$ and the parameter center $\boldsymbol{C}\in\mathbb{R}^{k\times n}$ can be represented the $\boldsymbol{A}\in\mathbb{R}^{m\times k}$ and $\boldsymbol{B}\in\mathbb{R}^{k\times n}$ in the original LoRA method. By these settings, we provide a better reparameterization of the tunable parameters in LoRA. It separates the parameters into different clusters, which have different functions in the whole network.
@@
After obtaining the matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, we further impose constraints on their product $\boldsymbol{AB}$ to align with the parameter distribution of the original weight matrix $\boldsymbol{W}_{0}$ as follows:
(12) $\min{\|{\boldsymbol{W_{0}}-\boldsymbol{AB}}\|_{2}}.$
@@
In this manner, we can ensure the initialization of $\boldsymbol{AB}$ is close to $\boldsymbol{W_{0}}$ at the beginning of the training.
@@
During the process of model adversarial adaptation, we design an adaptive update parameter, $\alpha$, to facilitate the model’s adaptive learning of robust semantic representations as follows:
(13) $\boldsymbol{Y}=\boldsymbol{XW_{0}}+\alpha\cdot\boldsymbol{XAB}.$
@@
$\alpha$ is a tunable neural network parameter, which can control the adaptation rate during the adversarial adaptation. In summary, we delineate the entire workflow of AdvLoRA in Algorithm 1.
@@
1:Images: $\boldsymbol{V}=\{{v}_{1},{v}_{2},\ldots,{v}_{n}\}$; Texts: $\boldsymbol{W}=\{{w}_{1},{w}_{2},\ldots,{w}_{n}\}$; Visual encoder: $\mathcal{F}_{v}$; Textual encoder: $\mathcal{F}_{w}$; Pre-trained weight matrix: $\boldsymbol{W}_{0}$; LoRA matrix: $\boldsymbol{A,B}$; Adaptive parameter: $\alpha$; Restriction value: $\epsilon$; PGD step: $\xi$; Loss function: $\mathcal{L}$.
2:Representations of $\boldsymbol{V}$ and $\boldsymbol{W}$: $\boldsymbol{Z^{v}}=\{\boldsymbol{z}_{1}^{v},\boldsymbol{z}_{2}^{v},\ldots,\boldsymbol{z}_{n}^{v}\}$, $\boldsymbol{Z^{w}}=\{\boldsymbol{z}_{1}^{w},\boldsymbol{z}_{2}^{w},\ldots,\boldsymbol{z}_{n}^{w}\}$.
3:while at adversarial fine-tuning stage do
4:    Perform clustering algorithm on $\boldsymbol{W_{0}}$ and obtain cluster center representation in Eq. (8) and Eq. (9);
5:    Obtain the LoRA matrix $\boldsymbol{A,B}$ from cluster center representation and $W_{0}$ in Eq. (10) and Eq. (11);
6:    Impose constraints on $\boldsymbol{A}$ and $\boldsymbol{B}$ with SGD algorithm in Eq. (12);
7:    Calculate the loss $l$ using $\boldsymbol{V}$, $\boldsymbol{W}$, $\boldsymbol{Y}$, $\mathcal{F}_{v}$, $\mathcal{F}_{w}$, and the loss function $\mathcal{L}$ in Eq. (5).
8:    Obtain the adversarial attack $\delta$ with $\boldsymbol{V}$, $\epsilon$, $k$, $\xi$ and $l$ in Eq. (3).
9:    Add $\delta$ to original images $\boldsymbol{V}$ to obtain the attacked images $\boldsymbol{V_{a}}$.
10:    Update $\boldsymbol{A},\boldsymbol{B}$ via Eq. (4).
11:end while
12:Generate robust representations of $\boldsymbol{V}$ and $\boldsymbol{W}$ to downstream tasks with adversarially adapted $\boldsymbol{A,B}$ and $\mathcal{F}_{v}$, $\mathcal{F}_{w}$.
@@ 4.1. Experimental Setup
Datasets. We conducted a comprehensive evaluation of our proposed model, AdvLoRA, across two types of retrieval tasks and four widely used datasets. This evaluation highlights AdvLoRA’s superior performance in cross-modal understanding tasks, specifically image-text retrieval with Flickr30K (Plummer et al., 2015) and MSCOCO (Lin et al., 2014), as well as video-text retrieval with DiDeMo (Anne Hendricks et al., 2017) and MSRVTT (Xu et al., 2016).
Baselines. We compare AdvLoRA with conventional adaptation methods, which are implemented by BLIP: full fine-tuning (BLIP-FFT), linear probe (BLIP-LP); as well as the PEFT method on BLIP: LoRA(BLIP-LoRA), Aurora, and Uniadapter.
  • BLIP-FFT (Li et al., 2022) is a conventional adaptation technique that enhances the performance of BLIP for specific downstream tasks by retraining and updating full parameters in downstream tasks.
  • BLIP-LP (Li et al., 2022) is an adaptation technique that involves adding and training a linear layer on top of the frozen pre-trained model BLIP to adapt to specific tasks.
  • BLIP-LoRA (Hu et al., 2021) is a PEFT technology that adapts BLIP by introducing low-rank adapters to capture task-specific information, allowing for efficient adaptation to downstream tasks with minimal tunable parameter updates.
Metrics. We employ $Recall@k$ as our evaluation metric, measuring the proportion of relevant items retrieved within the top $k$ results. A higher $Recall@k$ indicates better performance in retrieving relevant items, reflecting the model’s effectiveness and reliability in information retrieval tasks.
Implementations. Our implementation is based on Salesforce’s open-source codebase (Li et al., 2022). Following (Lu et al., 2023; Wang et al., 2023), we also apply BLIP (Li et al., 2022) as our vision-language backbone for all tasks. We use PyTorch to implement all experiments on the NVIDIA V100 GPU (32GB). We employ PGD-3 (Madry, 2017) for adversarial adaptation and to assess the model’s robustness. For the video-text retrieval task, we follow the work of Wei et al. (Wei et al., 2019) by adopting an attack strategy that sparsely samples 20% of the video frames. Furthermore, we adopt the setup of BLIP, utilizing a momentum encoder to enhance the retrieval performance of our model. To ensure a fair comparison, the momentum encoder is also applied to the other baseline methods. During the fine-tuning process, the parameters of the backbone model are kept frozen. We present the hyperparameter setting in Tab. 1.
Table 1. Hyperparameter setting
config Image-text Retrieval Video-text Retrieval
Flickr30K MSCOCO Didemo MSR-VTT
optimizer AdamW AdamW AdamW AdamW
lr 1e-5 1e-5 1e-4 1e-4
schedule cosine decay cosine decay cosine decay cosine decay
training bs 16 16 8 8
inference bs 32 32 8 8
frames - - 16 16
attack ratio - - 20% 20%
epochs 5 5 5 5
training input 384 384 8*224 8*224
inference input 384 384 16*224 16*224
attack type PGD-3 PGD-3 PGD-3 PGD-3
attack alpha 1/255 1/255 1/255 1/255
PGD-epsilon 1/255 1/255 1/255 1/255
rank 10 10 10 10
adaptive weight 1e-3 1e-3 1e-3 1e-3
weight norm lr 1e-3 1e-3 1e-3 1e-3
@@
Table 3. Adversarial experiment on Flickr30K. An asterisk (*) indicates that adversarial adaptation has been performed. The best results are displayed in bold, while the second-best results are underlined.